OpenStudy (anonymous):

Hey guys. I have a math quiz tomorrow on a concept I learned today in math, and there are a few problems that I need help with. She didn't fully explain how to do it, so I will be posting them in a second. The topic is completing the square.

OpenStudy (anonymous):

1200+9X=25X OR 1200=25X-9X OR 1200=16X OR X=1200/16 OR X=75 BAGS TO BREAK EVEN. PROOF 1200+9*75=25*75 OR 1200+675=1875 OR 1875=1875 sooo yea medal?????

OpenStudy (anonymous):

oops wrong problem

OpenStudy (anonymous):

\[9x ^{2}+12x+4=5\]